19. Which of the following is an even function? 1) f(x) = 3x2 2) f(x) = 3 + x 3) f(x) = sin(3x) 4) f(x) = log10 3x
Answer #1

Solve Even functions f(-x) = f(x) 0 Flx) = 382 F(-X)- 37-x)2 = 3x2 -FIX) = f(x)=3y² b even function flaga 3+X fl*) 3.2 3 PIX)